Mathematics 134 – Intensive Calculus for Science 2 Lab Pair Assignments January 19-20, 2006

Group 1: Group 11: Almeida, Carla Singleton,Paige Bochacki,Caelainn Doig, Pam

Group 2: Group 12: Brayton,Ali Allen, Tim Cherackal,Jo-Jo Van Dongen, Jennifer

Group 3:

Descombes,Samanta Dieguez,Felipe

Group 4:

Doan,Jimmy Turkson,Francisca

Group 5:

Donohue,Kate Everhart,Andrea

Group 6:

Flaherty,Elisabeth Flood,Brian

Group 7:

Jones,Melissa Lanzillo,Lauren

Group 8:

Leon Hereter,Camille Linga,Rebekah

Group 9:

Marino,John McNaught,Elizabeth

Group 10:

Moss,Sam Panko,Laura